OpenStudy (anonymous):

HELP Factor the expression completely. Begin by factoring out the lowest power of each common factor. 12x^(−1/2) + 7x^(1/2) + x^(3/2)

OpenStudy (anonymous):

use wolfram Alpha

jhonyy9 (jhonyy9):

12/radical x + 7 radical x + x radical x = 12/radical x + (7+x)radical x = =(12+x(7+x))/radical x